OpenStudy (naudiapie):

The pilot of a rescue helicopter is flying over the ocean at an altitude of 1250 ft. The pilot sees a life raft at an angle of depression of 31°. The horizontal distance from the helicopter to the life raft, rounded to the nearest foot is ____ feet. (Enter only the number.)

OpenStudy (naudiapie):

?????

OpenStudy (naudiapie):

pls help

OpenStudy (naudiapie):

@pooja195

OpenStudy (naudiapie):

@Algorithmic

OpenStudy (naudiapie):

@perfectdawn

OpenStudy (anonymous):

 Are you familiar with SOH CAH TOA at all? It will really help here.

OpenStudy (naudiapie):

not really

OpenStudy (naudiapie):

well i learned it but i forget

OpenStudy (anonymous):

Okay, that fine. SOH goes with Sine which is Opposite/Hypotenuse CAH goes with Cosine which is Adjacent/Hypotenuse TOA goes with Tangent which is Opposite/Adjacent In my previous post I added a picture to help. By looking at the angle given, we can see that we are looking at the opposite and adjacent sides, since we aren't using the hypotenuse. This means we'll be using TOA for our problem. Any questions so far? I know triangles can be a bit tricky.

OpenStudy (anonymous):

Normally we do \[\tan (x)=\frac{ opposite }{ adjacent }\] but since we are given the angle needed, we have to do \[\tan^{-1} (31)=\frac{ opposite }{ adjacent }\] As we keep filling in missing pieces, we see the opposite is what we need to find and we know the adjacent side is 1250.\[\tan^{-1} (31)=\frac{ x }{ 1250 }\] Now it's just simple multiplication by 1250 on both sides.\[\tan^{-1} (31)\times1250=x\] You'll need a calculator to solve this now, but your answer should be the horizontal distance between the helicopter and the life vest. Hope this helped!